OpenStudy (shoba):

2 HOCl + 2 C2H4O2 = Cl2 + 2 C2H3O2 + 2 H2O from these equation how to calculate the mass of HOCl and C2H4O2 to produce 5ppm of chlorine gas?

OpenStudy (nincompoop):

yeah, first make sure that the reaction is balanced

OpenStudy (nincompoop):

then get your periodic table of elements, because you will be calculating moles and converting them into grams

OpenStudy (nincompoop):

now if that reaction is already balanced (something you will have to make sure) you are looking at Cl2 and it tells you that for the reactants on the left hand side yields 1 mole of Cl2 make sure you know the mass of Cl2

OpenStudy (shoba):

the equation is balanced.

OpenStudy (nincompoop):

how many grams are there in a mole of Cl2

OpenStudy (shoba):

70.906 gram

OpenStudy (nincompoop):

how many ppm is that?

OpenStudy (nincompoop):

1 ppm = 0.001 grams

OpenStudy (shoba):

i tell based on the molar mass actually....if 70.906 gram means it will be 70906 ppm?

OpenStudy (nincompoop):

yeah because you move the decimal point to the right three times

OpenStudy (nincompoop):

you only need 5, so that means you need a smaller amount of reactants

OpenStudy (nincompoop):

WHAT?

OpenStudy (shoba):

sorry how u know 1ppm = 0.001g?

OpenStudy (nincompoop):

my bad, I was thinking of something else 0.001g/litre = 1 ppm

OpenStudy (shoba):

so 1mg/l = 1ppm am i right? so how to solve this??

OpenStudy (nincompoop):

dynamic conversion

OpenStudy (shoba):

dynamic conversion means?

OpenStudy (shoba):

i need to find the mass of hocl and c2h4o2 to produce 5 ppm of cl gas.....

OpenStudy (nincompoop):

\[5 \times (\frac{0.001g}{1L}) = 5ppm\]

OpenStudy (nincompoop):

you have to look at the reaction and reactants your reactants yield 1mole of Cl2

OpenStudy (shoba):

yes two mole of hocl and c2h4o2 produce one mole of cl2..but how to calculate..:(

OpenStudy (nincompoop):

have you learned this part?

OpenStudy (nincompoop):

your reactants are 2:2 so that means it is a 1:1 ratio

OpenStudy (shoba):

i dont actually

OpenStudy (nincompoop):

a 1:1 ratio is easy to work with just like when you need 1 table spoon of sugar for every 1 table spoon of flour when baking that yields 2 table spoon of doughnut and if you want 5 table spoon of doughnut you will have to increase your sugar and flour your ratio starts with 1:1:2 now you want x:y:5 first, (1:1:2)/2 = 0.5:0.5:1 since we want 5, we multiply by 5 5(0.5:0.5:1) = (2.5:2.5:5)

OpenStudy (nincompoop):

now with what you have, you have more than 1 reactions but you only need to worry about the Cl2 so with 2:2:1 all in moles 1 mole of Cl2 = 70.906g of Cl2 70.906 g Cl2 = 70906mg of Cl2 you just need 5.0 grams of Cl2 in a litre to make 5 ppm (70906/1L) but needs 5/1L since they both need to be in Litre anyway, we can drop that for now 70906x = 5 solve for x remember that you don't yield Cl, but instead you are yielding Cl2 and that is something you need to account for Cl2 -> Cl+Cl

OpenStudy (nincompoop):

now a more systematic approach would be, instead of using moles, use grams. convert the moles of the reactants into grams A:B:C you already know C (assuming that you will use that as your Cl2)

OpenStudy (nincompoop):

2:2:1 2A:2B:1C 1A:1B:0.5C 0.5A:0.5B:0.25C are you getting this ratio technique?

OpenStudy (shoba):

yes i get this........

OpenStudy (nincompoop):

sorry how many moles of Cl2 do you need to get 5mg of Cl2?

OpenStudy (shoba):

0.07 mole of chlorine

OpenStudy (shoba):

thanks alot @nincompop i able to solve this question already...... :) 2 HOCl + 2 C2H4O2 = Cl2 + 2 C2H3O2 + 2 H2O in order to produce 5 ppm of chlorine gas we need to the moles first... 5ppm = (5mg/L)*(1mole/7.09mg) = 0.0705mol/L from the ratio theoram 2:2:1=(0.141:0.141:0.0705) now we find the mass of HOCL = 0.141*52.4603 = 7.397 mg/l c2h4o2 = 0.141*60.05196 = 8.47mg/l